Hartford Research issues bonds dated January 1, 2017, that pay interest semiannually on June 30 and December 31. The bonds have a $40,000 par value and an annual contract rate of 10%, and they mature in 10 years. (Table B.1, Table B.2, Table B.3, and Table B.4) (Use appropriate factor(s) from the tables provided. Round all table values to 4 decimal places, and use the rounded table values in calculations. Round your 'Present Value' answers to the nearest whole dollar.)
  
Required:
Consider each of the following three separate situations.
  
1. The market rate at the date of issuance is 8%.
(a) Complete the below table to determine the bonds' issue price on January 1, 2017.
(b) Prepare the journal entry to record their issuance.
2. The market rate at the date of issuance is 10%.
(a) Complete the below table to determine the bonds' issue price on January 1, 2017.
(b) Prepare the journal entry to record their issuance.
3. The market rate at the date of issuance is 12%.
(a) Complete the below table to determine the bonds' issue price on January 1, 2017.
(b) Prepare the journal entry to record their issuance.

Complete the below table to determine the bonds' issue price on January 1, 2017, if the market rate at the date of issuance is 8%.

Table values are based on:
n =
i =
Cash Flow Table Value Amount Present Value
Par (maturity) value
Interest (annuity)
Price of bonds

Journal Entry: Record the issue of bonds with a par value of $40,000 on January 1, 2017. Assume that the market rate of interest at the date of issue is 8%.

Complete the below table to determine the bonds' issue price on January 1, 2017, if the market rate at the date of issuance is 10%.

Table values are based on:
n =
i =
Cash Flow Table Value Amount Present Value
Par (maturity) value
Interest (annuity)
Price of bonds

Journal Entry: Record the issue of bonds with a par value of $40,000 on January 1, 2017. Assume that the market rate of interest at the date of issue is 10%.

Complete the below table to determine the bonds' issue price on January 1, 2017, if the market rate at the date of issuance is 12%.

Table values are based on:
n =
i =
Cash Flow Table Value Amount Present Value
Par (maturity) value
Interest (annuity)
Price of bonds

Journal Entry: Record the issue of bonds with a par value of $40,000 on January 1, 2017. Assume that the market rate of interest at the date of issue is 12%.
Answer #1
Concepts and reason

Bonds payable: It is an instrument issued by an entity which may be either corporate or government for raising fund by paying fixed or variable interest for the term of the bond. It can either be issued at discount or at premium.

Fundamentals

Journal entry: Journal entry means recording of a transaction which are related to the business of the company in the records of the company.

Present value: Sum of money which is receivable or payable at a future date is discounted to arrive at current value. This value is the present value of future sum of money.

Present value of annuity(PVA)
: Sum of money which is receivable or payable at a different future dates is discounted to arrive at current value. This value is the present value of future value of annuity.

Table value for maturity value = .
(1+i) 20
(1+49
20 =0.4564

Table value for interest = =
(1+i)?
(1+i)
Tube value for more
- (14%) *(1 +4ya t..+(14%** *(1+4)
= 13.5903

Premium on bonds payable = Issue price - Maturity value
= $45,437 - $40,000
= $5,437

1
Table value for maturity value = .
(1+i)
(1+5% 20 = 0.3769

for interest = () (hiyo.* (1* *(1+)
= (1+5%) * (1+5%) _* (1459.)*** (15%)
= 12.4622

There will be no discount or premium on bonds payable when market rate is 10%.

Table value for maturity value = .
(1+i)
(1+692 20 = 0.3118

Tube value for more
Table value for interest = =
(1+i)?
(1+i)
(10* (-4023 * mm (100
km
(1+6%)
=11.4699
(1+6%)

Discount on bonds payable = Maturity value - Issue price
= $40,000-$35,412
= $4,588

Ans: Part 1a

Table values are based on:
20
Cash flow
Par (maturity) value
Interest (annuity)
Price of bonds
Table value Amount
0.4564 40,0

Part 1b

Credit ($)
Debit ($)
$45,437
Date
Accounts Title and Explanation
1-Jan-17 Cash
Bonds Payable
Premium on Bonds Payable
|(to re

Part 2a

Table values are based on:
20
5
Cash flow
Par (maturity) value
Interest (annuity)
Price of bonds
Table value Amount
0.3769 40

Part 2b

Credit ($)
Debit ($)
$40,000
Date
Accounts Title and Explanation
1-Jan-17 Cash
Bonds Payable
|(to record issue of bonds)
$40,

Part 3a

Table values are based on:
201
6
Cash flow
Par (maturity) value
Interest (annuity)
Price of bonds
Table value Amount
0.3118 4

Part 3b

Credit ($)
Date
Accounts Title and Explanation
1-Jan-17 Cash
Discount on Bonds Payable
Bonds Payable
|(to record issue of bon